OpenStudy (anonymous):

Find the number of real solutions depending on values of k in 8^[x(x+k)] = [1/16]^(27/4)

OpenStudy (anonymous):

I got it down to the equation x^2 + kx + 9 = 0

jimthompson5910 (jim_thompson5910):

8^[x(x+k)] = [1/16]^(27/4) (2^3)^[x(x+k)] = [2^(-4)]^(27/4) 2^[3x(x+k)] = 2^(-4*27/4) 3x(x+k) = -27 3x^2 + 3k = -27 3x^2 + 3k + 27 = 0 The last equation above will have real solutions when \(\Large D \ge 0\) D is the discriminant is given by D = b^2 - 4ac So 3x^2 + 3k + 27 = 0 will have real solutions when \(\Large b^2 - 4ac \ge 0\) In this case, a = 3, b = 3 and c = 27.

jimthompson5910 (jim_thompson5910):

oops, i mean a =3, b = 3k, and c = 27 so \(\Large b^2 - 4ac \ge 0\) \(\Large (3k)^2 - 4(3)(27) \ge 0\) keep going to solve for k

OpenStudy (anonymous):

I know you get k^2 > 36

jimthompson5910 (jim_thompson5910):

good, you're very close

OpenStudy (anonymous):

\[\left| k \right| > 6\]

jimthompson5910 (jim_thompson5910):

you can write the answer as \[\Large |k| \ge 6\] or as \[\Large k \le -6 \ \text{or} \ k \ge 6\]

OpenStudy (anonymous):

But the question isn't asking me to solve for k. It says "find the number of real solutions depending on values of k"

jimthompson5910 (jim_thompson5910):

Does it give you values of k to use?

OpenStudy (anonymous):

no. That's why I was slightly confused. I heard from a friend that the answer is supposed to be 2 real values. But I'm not sure why.

jimthompson5910 (jim_thompson5910):

Maybe there's more context elsewhere?

OpenStudy (anonymous):

Oh wait I think I have an idea:

OpenStudy (anonymous):

So above the part of the sheet we got with the questions, there is some information- it talks about the discriminant, and how the values of the discriminant affect how many roots there are and the nature of the roots it says that if b^2 - 4ac > 0, there are two real roots. So would the answer be that if k≤−6 or k≥6, then there are two real roots?

jimthompson5910 (jim_thompson5910):

if you mean two distinct real roots, then yes b^2 - 4ac > 0 which means that if k < -6 or k > 6 then the original equation has 2 distinct real roots

OpenStudy (anonymous):

i think that's what she wants. can you help me with another one?

OpenStudy (anonymous):

(and thank you)

jimthompson5910 (jim_thompson5910):

sure, what is it

OpenStudy (anonymous):

should i post it within this question?

jimthompson5910 (jim_thompson5910):

either way works

OpenStudy (anonymous):

Find all values of k for which roots of the equation: \[5^{x(x+1)}\times25^{[k(k-1)]/[2]} = \sqrt{5^{x ^{2}}}\times125^{(kx+k+1)/(2)}\] satisfy the restriction: 1/x1 + 1/x2 > 0

OpenStudy (anonymous):

(satisfy the restriction: \[1/x _{1} + 1/x _{2} > 0\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{x_{1}}+\frac{1}{x_{2}} > 0\] \[\Large \frac{x_{2}}{x_{1}x_{2}}+\frac{x_{1}}{x_{1}x_{2}} > 0\] \[\Large \frac{x_{2}+x_{1}}{x_{1}x_{2}} > 0\] \[\Large \frac{x_{1}+x_{2}}{x_{1}x_{2}} > 0\] So \[\Large \frac{1}{x_{1}}+\frac{1}{x_{2}} > 0\] is equivalent to \[\Large \frac{x_{1}+x_{2}}{x_{1}x_{2}} > 0\]

jimthompson5910 (jim_thompson5910):

5^[x(x+1)]*25^[k(k-1)/2] = sqrt(5^(x^2))*125^[(kx+k+1)/(2)] 5^[x(x+1)]*(5^2)^[k(k-1)/2] = (5^(x^2))^(1/2)*(5^3)^[(kx+k+1)/(2)] 5^[x(x+1)]*5^[2k(k-1)/2] = 5^((1/2)*x^2)*5^[3(kx+k+1)/(2)] 5^[x(x+1)+2k(k-1)/2] = 5^((1/2)*x^2)+3(kx+k+1)/(2)] x(x+1)+2k(k-1)/2 = (1/2)*x^2+3(kx+k+1)/(2) x^2 + x + k^2 - k = (1/2)*x^2 + (3kx + 3k + 3)/2 2x^2 + 2x + 2k^2 - 2k = x^2 + 3kx + 3k + 3 2x^2 + 2x + 2k^2 - 2k - x^2 - 3kx - 3k - 3 = 0 x^2 + (2 - 3k)x + (2k^2 - 5k - 3) = 0 In this case, a = 1, b = 2 - 3k, and c = 2k^2 - 5k - 3 Solve for x. x = (-b+-sqrt(b^2-4ac))/(2a) x = (-(2 - 3k)+-sqrt((2 - 3k)^2-4*1*(2k^2 - 5k - 3)))/(2*1) x = (-2 + 3k+-sqrt(4 - 12k + 9k^2-8k^2 + 20k + 12))/(2*1) x = (-2 + 3k+-sqrt(k^2 + 8k + 16))/(2) x = (-2 + 3k+-sqrt((k+4)^2))/(2) x = (-2 + 3k+-|k+4|)/(2) x = (-2 + 3k+|k+4|)/(2) or x = (-2 + 3k - |k+4|)/(2) So the two solutions for the original equation are x = (-2 + 3k+|k+4|)/(2) or x = (-2 + 3k - |k+4|)/(2) Since we have a +/-, this takes care of the absolute value, so we can say that the two solutions are x = (-2 + 3k+(k+4))/(2) or x = (-2 + 3k - (k+4))/(2) x = (4k+2)/(2) or x = (2k - 6)/(2) x = 2k + 1 or x = k - 3 So we can say x1 = 2k + 1, x = k - 3 But above we know that (x1 + x2)/(x1*x2) > 0 So this really means (2k+1 + k -3)/((2k+1)(k-3)) > 0 (3k-2)/((2k+1)(k-3)) > 0 Keep going to solve for k

OpenStudy (anonymous):

so i have [3k - 2] / [(2k+1)(k-3)] > 0

OpenStudy (anonymous):

I'm actually not sure what to do from there

OpenStudy (anonymous):

I can't multiply both sides by 2k^2 - 5k - 3 because I could be multiplying by zero

jimthompson5910 (jim_thompson5910):

look for the zeros of [3k - 2] / [(2k+1)(k-3)] also, look for the vertical asympotes of [3k - 2] / [(2k+1)(k-3)]

OpenStudy (anonymous):

Are you telling me to graph it...?

jimthompson5910 (jim_thompson5910):

you can if you want, but you can also determine this info without graphing

OpenStudy (anonymous):

How? (also I haven't learned asymptotes yet)

jimthompson5910 (jim_thompson5910):

find the values of k which make the denominator zero

OpenStudy (anonymous):

k = -1/2 or k = 3

OpenStudy (anonymous):

What do I do with those values? The denominator can't be zero

jimthompson5910 (jim_thompson5910):

3k - 2 = 0 k = 2/3 So we must take this value into consideration as well

jimthompson5910 (jim_thompson5910):

so set up a number line with the values, k = -1/2, k = 3 and k = 2/3 on it

jimthompson5910 (jim_thompson5910):

then test values between these numbers to see where [3k - 2] / [(2k+1)(k-3)] is positive

OpenStudy (anonymous):

why only between?

jimthompson5910 (jim_thompson5910):

because the actual values themselves either give 0 or give you a division by zero error

OpenStudy (anonymous):

but can't you go past the values? for example, test k = 9 or test k = -7

jimthompson5910 (jim_thompson5910):

you can test as many values as you want, but it only makes sense to test values from each region on the number line

OpenStudy (anonymous):

okay so -1/4 works, 1/3 works, and 1 does not work.

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

then what?

jimthompson5910 (jim_thompson5910):

these values represent intervals, you're missing one interval though

OpenStudy (anonymous):

Which one? I have the interval between 0 and -1/2, 0 and 2/3, and 2/3 and 3

jimthompson5910 (jim_thompson5910):

how about from 3 to infinity?

OpenStudy (anonymous):

wait but i asked that. i asked if you should go beyond the limits of -1/2 and 3 and you said no. then do i also need to test something between -1/2 and - infinity?

OpenStudy (anonymous):

("but can't you go past the values? for example, test k = 9 or test k = -7 15 minutes ago"

jimthompson5910 (jim_thompson5910):

why not? they're test values. If they fail the test, then you ignore that interval

OpenStudy (anonymous):

-5 doesn't work but 6 works. so the ones that work are between -1/2 and 0, 0 and 2/3, and 3 and infinity

jimthompson5910 (jim_thompson5910):

good, those are the intervals where k makes [3k - 2] / [(2k+1)(k-3)] positive

OpenStudy (anonymous):

oh so is that the answer? how would i formate something like that?

OpenStudy (anonymous):

*format

jimthompson5910 (jim_thompson5910):

you can use interval notation, or say k > 3 for the interval from 3 to infinity for instance

OpenStudy (anonymous):

we haven't learned/worked with interval notation. could i say:

OpenStudy (anonymous):